OpenStudy (anonymous):

ellipse: (x^2)/25 + (y^2)/16 =1 find the endpoints of the major and minor axises, center, foci, & eccentricity

OpenStudy (anonymous):

major axis (+_5,0) minor axis (+-4,0) center (0,0) eccen.=sqrt(1-16/25)=3/5

OpenStudy (anonymous):

foci=(+-3,0)